What is the frequency of a photon with a wavelength of 159 nm. Report your answers to three significant digits.
1
Expert's answer
2020-10-10T08:36:03-0400

the frequency of a photon with a wavelength, λ = 159 nm

= 159×10-9 m

Let, frequency of the Photon = v s-1


Then, velocity of photon (light)

C = λ×v m.s-1

Or, v = C/λ s-1

As we know C =3×108 m.S-1 Or, v = 3×108/159×10-9

Or, v = 1.88 ×1015

(3 significant digits)


Hence, the frequency of the Photon = 1.88 ×1015 S-1